OpenStudy (anonymous):

okay so there were 30 questions on my math test and each math test is 100 point how many can I miss and still get an A?

OpenStudy (anonymous):

depends on what each question is worth some could be worth 5 and others could be worth 1.

OpenStudy (anonymous):

I think there all worth the same because it would have said if it was

OpenStudy (anonymous):

each question is worth 3.3 points. If you get 26 questions right (86%) you will have an A.

OpenStudy (anonymous):

If they were all the same point value

OpenStudy (anonymous):

\[\frac{ 30-3}{ 30 }=\frac{ 27 }{ 30 }=90%\]

OpenStudy (lukecrayonz):

What is the value of an A? 90 or 95?

OpenStudy (anonymous):

85% is an A where im from

OpenStudy (anonymous):

90 percent is a A 85 is a B. Where are from XD

OpenStudy (anonymous):

3 incorrect answers is the most you can get to keep an A.

OpenStudy (anonymous):

Canada :) i got it easy i guess.

OpenStudy (anonymous):

If an A is 90% or Above.

OpenStudy (anonymous):

well different parts of Canada are different..

OpenStudy (anonymous):

Haha you do! and i missed to already :(

OpenStudy (anonymous):

I mean 2 questions

OpenStudy (anonymous):

Hope you do well!

OpenStudy (anonymous):

You're still good though.

OpenStudy (anonymous):

hopefully but oh well